In the electrolysis of acidulated water, it is desired to obtain 1.12 cc of hydrogen per second under S.T.P. conditions. The current to be passed is
1.0.965 Amp
2.19.3 Amp
3.9.65 Amp
4.1.93 Amp

Grade:12

1 Answers

Vikas TU
3 years ago
No of moles of H2=1.12/22400
∴ No. of equivalents of hydrogen=1.12×2/22400=10^−4
No. of faradays required=10^−4
∴Current to be passed in 1 sec. =96500×10^−4
=9.65A